In 4 years time, Sam will be twice as old as he was 4 years ago. What is his current age?
Answer by macston(5194)

.
S=Sam's age
.
S+4yrs=2(S-4yrs)
S+4yrs=2S-8yrs
S+12yrs=2S
12yrs=S
.
ANSWER: Sam is 12 years old.